Shift report 20030507 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed May 07 2003 - 21:10:42 EDT

  • Next message: Apache: "Shift report 20030508 00:00-08:00"
    ------------------------------------------------------------
    Date: Wed May  7 21:10:42 2003, Shift: 16:00-24:00
    Supervisor: fv, with: 
    Log book#14 pages 94 - 94
    
    Shift Summary:
    ------------------------------------------------------------
    The shift was devoted to polarization development
    particular in yellow as apperently decided in th eWed coordination metting.
    SInce no beam is expected before midnight I will stop the shift now.
    
    a) Magnets left on standby.
    b) The H1 ch 1-16 bottom has a bad HV card .All voltages are
       up by ~ 20 V. This is most likely the reason for the     failure of the near online following run 8812. With a new timing cal it is again quite good.
    Cards should be fixed at next access.
    c) Cabled up a few more signals lines to polarization counters.
    
    d) The BFS front camera is dimmed (light/ video interface?)
      should be checked too.
    
    e) No more cookies !
    
    
    
    
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed May 07 2003 - 21:11:35 EDT